A scientist wants to make 6 milliliter of a 30% sulfuric acid solution. The solution is to be made from a combination of a 20% sulfuric acid solution and a 50% sulfuric acid solution. How many milliliters of each solution must be combined to make the 30% solution?
Answer by jorel1380(3719)

6x.3=1.8
m+n=6
.2m+.5n=1.8
2m+5n=18
2m+2n=12
3n=6
n=2
m=4
You need 4 ml of 20% solution and 2 ml of 50% solution to make 6 ml of 30% solution.